OpenStudy (anonymous):

Find all solutions to the following system of linear equations : x+y-2z=-9 3x-5y+2z=-11 4x-7y+3z=-14

OpenStudy (anonymous):

solve for one variable first

OpenStudy (anonymous):

or you can add 3x-5y+2z=-11 to x+y-2z=-9

OpenStudy (anonymous):

that should make it easier.

OpenStudy (anonymous):

21x+5y=94 ... 2x–3y+2z=2 x+ 4y-z=9 -3x+y–5z=5 so far, I have, 2 -3 2 |2 1 4 1 |9 -3 1 -5 |5 ... y-z=11 z=29 x=-122 y=40 z=29. I plugged in my solutions to the equation to ... method you would use to solve the following system of equations and why. ... use another method to determine the solution set. x - y + 2z = 6 4x + z = 1

OpenStudy (anonymous):

14) x - 2y + 2z = 3 ... 3) x + 2y + 3z = 9. 3x + 4y + z = 5. 2x – y + 2z = 11. 4) x + 2y = 0. y + z = 3. x + 3z = 14 ... 9) Find three solutions to the following system of equations. x + ... 12) Why is it not possible for a linear system to have exactly two solutions? ... 7) 3x – 5y = 2. –x + 2y = 0. 8) x + 2z = 8. y + 4z = 8. z = 3. 9) x + y – z = 2.

OpenStudy (anonymous):

Determine whether each ordered pair is a solution of the system of linear ... 4, 4 8, 0 1, 5 10, 4 1, 2 2. 3. 4. 0.5x 1.5y 7 x 2y 4 6x 9y 9 2x 6y 28 3x 2y 4 2x 3y 6 In ... the linear system. 11. 5, 2 8, 12. x 3x 3y y 6 8 1 3,1 4x x y 3y 9 16 13. 14. x 4z 17 ... 2 x 5y 3z 19 9a 3b c 4 In Exercises 15 and 16, write a system of linear equations ...

OpenStudy (anonymous):

c) 2x – 8y + 7z = 6 d) 4x + 2y – 3z + 5w = 9. e) 6x + 4y – 4z + 8w – 2t = 5 f) 7x1 + 5x2 – x3 + 2x4 – 9x5 + 3x6 + 4x7 + x8 = 1 ... 3) 8x + 2y – 3z + 4w = 4 4) 4x – 5y + 2z – 3w + 6t = 0. -x + .... It turns out that not all systems of linear equations have solutions. ... To find all solutions, all we need to do is solve for y to obtain y = 2x – 4.

OpenStudy (anonymous):

For each of the following pair of equations find the points of intersection: ... (e) -3x + 5y = 11, x = 5y - 14. (-3, 115). (f) 2x - y = 4, y = -3x - 9. (-1, -6). (g) 4x - 3y ... + 1, 2x - 2y = -2 The same line. All solutions to y = x + 1. (i) y = 2x + 3, y = 5x + 6. (-1, 1). 2. ... 7x - 3y + 2z = -25. -3x + 2y + 3z = 35 x + y + z = 10 x = -2,y = 7,z = 5.

OpenStudy (anonymous):

x+2y+3z=5 | x-y+6z=1 | 3x-2y=4 \ y+4z=8 / x+y+3z-4t=12 \ 3x+y-2z-t=0 / 2x+3y+4z=5 ... Therefore we can gather all essential data of the system in one matrix, by adding the ... Two systems are equivalent if and only if they have the same set of solutions. .... x-3y=21 | 4x+2y=14 \ 3x+3y=7 [1 -3 21] [4 2 14] [3 3 7] (1/2).R2 ; R3-3.

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (anonymous):

wow...did you get your answers from that or not ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok...just checking :)

OpenStudy (the_fizicx99):

Lol speed down through all of that and its looks Chinese >.e

OpenStudy (anonymous):

I know...I would have just used elimination

OpenStudy (anonymous):

thanks